# CAVIAR: Categorical-Variable Embeddings for Accurate and Robust Inference

### 07 Apr 2024 v1

@@ Introduction
Given there are only 41,700 zip codes in the United States, it is conceivable that at some point, the zip code variable would saturate, and additional data would only accrue for zip codes that were already included in the model by virtue of earlier observations. However, in most practical data samples, especially those with fewer than 41,700 observations, this saturation is unlikely to occur.
@@
While the coding schema for zip codes in this example is well-defined and finite, the complexity of real-world data often presents scenarios with no upper limit to the number of categorical levels. For instance, consider including ‘brand’ as a categorical variable in our merchant example. As the product line evolves over time, with new brands being introduced and older brands becoming dormant, the number of levels in this variable ($L_{n}$) will inherently increase with the sample size ($n$) and become potentially unbounded. This phenomenon is not limited to brands alone; similar observations can be made for store locations, product attributes, and other variables—all of which may vary over time such that new values emerge and old values become irrelevant. As such, any dynamic element in the research environment that is described using a categorical variable will inevitably lead to a dynamically varying definition of categorical variable levels as new levels are constructed to track emergent aspects and extant levels become dormant with environmental changes.
@@
Moreover, behind these assertions lies the expectation that $\beta$ is crucial to the research question. To further develop this example, consider a scenario where a firm seeks to determine where it should focus its marketing efforts. Using zip code as a categorical variable could serve as a natural proxy to understand variations in factors such as demand across different zip codes. In such cases, accurately estimating $\beta$ while maintaining the granularity of the corresponding categorical variable may be essential to the analysis. However, while researchers may consider devising alternate schemas—such as merging ‘neighboring’ category levels like adjacent zip codes—such approaches would inevitably lead to a loss in the specificity of findings; a dance between specificity and estimation robustness is foundational to the specification of a categorical schema in the econometric model. Additional examples may include the Fama-French model, where the firm-specific intercepts (designated as ‘alpha’ in the canonical treatment) may often be more critical than the coefficients on the four factors. [^1] A firm-specific alpha (i.e., $\alpha_{f}$ in the Fama-French literature) is represented as $\alpha+\beta_{f}$ in our specification, where $\beta_{f}$ denotes the deviation of firm $f$’s coefficient from the grand mean.
@@
As panel fixed effects are not pivotal, a common route to estimation is differencing sequential observations to remove the influence of the fixed effect—an approach that yields consistent population-level parameter estimates at the expense of being uninformative about the fixed effects (Arellano and Bond 1991). This solution, however, is less desirable when the manifold can be formed through structured and unstructured variables such that the categories are on or near it, as then the categorical variable fixed effects can be recovered in addition to the population parameters. This aspect is the focus of our contributions.
Last, we note that the partially linear setup is typical in high-dimensional categorical variables. More sophisticated variants, such as the inclusion of interactions between $f(x_{i};\zeta)$ and $D_{li}$, can be considered in our framework and would likely yield model structures that also hew close to established econometric models. However, the identification challenge stems from the high dimensionality of $\{\beta_{l}\}_{l=1}^{L_{n}}$. Our approach innovates by providing a higher-order model structure to reduce data demands and drive identification. While such structures can be ported to more complex cases, this would come at the cost of increased data demands that may be simpler to list in theory than to satisfy in practice; depending on the dimensionality of the categorical variable and the underlying manifold, more complex models may or may not be practicable even with our innovations.
@@
Returning to our merchant example, consider embedding zip codes in a 2-dimensional space where the axes correspond to longitude and latitude. This approach may provide valuable insights. For example, if a retailer sells apparel suited to specific temperatures (e.g., T-shirts), projecting the high-dimensional categorical variable onto a 2-dimensional space and including each corresponding location as data features could offer a nuanced understanding of the economic problem. The model could be easily expanded by incorporating additional features, such as elevation. Imagine, in addition to elevation, we have data on attributes like rainfall. Suppose we augmented the geographic variables with a large language model (LLM) embedding of the verbal description of each zip code. These data features may prove useful as they encompass not only the geographic characteristics of each zip code but also information on economics, culture, etc.
@@ Conclusion
Our method has broad applicability across various research fields and industry sectors where high-dimensional categorical data is prevalent. For example, in healthcare analytics, patient data often includes categorical variables such as diagnosis codes and treatment codes. Extending our approach could involve using the LLM encoding of each diagnosis code or treatment code as a set of explanatory covariates for the fixed effect. This would transform the high-dimensionality non-parametric model, which might result from including many applicable diagnosis and treatment codes, into a lower-dimensional, computationally stable, and tractable model. Such approaches may be particularly beneficial if data describing the category levels is not only structured or textual but also multi-modal, allowing for a lower-dimensional embedding to be derived from visual images, among other sources. For instance, images of various diagnoses (e.g., an X-ray image of a fracture) could serve as inputs to the embedding to bolster its predictive accuracy. We hope our method and findings prove useful in such endeavors.

### 11 Apr 2024 v2

@@ Introduction
Or, in education sciences, consider assessments of educational efficacy in school districts or teacher programs. A longitudinal study tracking student outcomes across multiple cohorts may start with a limited number of school districts but gradually expand to include more districts as the study progresses. As more data is collected over time, the number of distinct school districts or teacher education programs represented in the data may increase (Darling-Hammond et al. 2005, Hanushek 2016). In a categorical variable representing school districts, the inclusion of these additional category levels, driven by the growing scope of the data, would lead to the same challenges of high dimensionality and sparsity.
@@
Given there are only a finite number of zip codes or school districts in the United States, it is conceivable that at some point, a categorical variable of zip codes or school districts would saturate, and additional data would only accrue for zip codes and school districts that were already included in the model by virtue of earlier observations. However, in most practical data samples, for instance those with fewer than 41,700 observations in data where there are 41,700 zip codes and all zip codes are candidates for inclusion, saturation will not occur.
@@
Moreover, while the coding schema for zip codes and school districts is well-defined and finite, the complexity of real-world data often presents scenarios with no upper limit to the number of categorical levels. For instance, consider including ‘brand’ as a categorical variable in our merchant example. As the product line evolves over time, with new brands being introduced and older brands becoming dormant, the number of levels in this variable ($L_{n}$) will inherently increase with the sample size ($n$) and become potentially unbounded. This phenomenon is not limited to brands alone; similar observations can be made for store locations, product attributes, and other variables—all of which may vary over time such that new values emerge and old values become irrelevant. As such, any dynamic element in the research environment that is described using a categorical variable will inevitably lead to a dynamically varying definition of categorical variable levels as new levels are constructed to track emergent aspects and extant levels become dormant with environmental changes.
@@
Behind these assertions lies the expectation that $\beta$ is crucial to the research question. To further develop this example, consider a scenario where a firm seeks to determine where it should focus its marketing efforts. Using zip code as a categorical variable could serve as a natural proxy to understand variations in factors such as demand across different zip codes. In such cases, accurately estimating $\beta$ while maintaining the granularity of the corresponding categorical variable may be essential to the analysis. However, while researchers may consider devising alternate schemas—such as merging ‘neighboring’ category levels like adjacent zip codes—such approaches would inevitably lead to a loss in the specificity of findings; a dance between specificity and estimation robustness is foundational to the specification of a categorical schema in the econometric model. Additional examples may include the Fama-French model, where the firm-specific intercepts (designated as ‘alpha’ in the canonical treatment) may often be more critical than the coefficients on the four factors. [^1] A firm-specific alpha (i.e., $\alpha_{f}$ in the Fama-French literature) is represented as $\alpha+\beta_{f}$ in our specification, where $\beta_{f}$ denotes the deviation of firm $f$’s coefficient from the grand mean.
@@
As panel fixed effects are not pivotal, a common route to estimation is differencing sequential observations to remove the influence of the fixed effect—an approach that yields consistent population-level parameter estimates at the expense of being uninformative about the fixed effects (Arellano and Bond 1991). For instance, similar methodology has been used to estimate the effects of psychosocial job stressors on mental health while accounting for individual-specific factors such as some individuals being more likely to report better mental health than others (Milner et al. 2016). This solution, however, is less desirable when the manifold can be formed through structured and unstructured variables such that the categories are on or near it, as then the categorical variable fixed effects can be recovered in addition to the population parameters.
A partially linear setup is typical in high-dimensional categorical variables. More sophisticated variants, such as the inclusion of interactions between $f(x_{i};\zeta)$ and $D_{li}$, can be considered in our framework and would likely yield model structures that also hew close to established econometric models. However, the identification challenge stems from the high dimensionality of $\{\beta_{l}\}_{l=1}^{L_{n}}$. Our approach innovates by providing a higher-order model structure to reduce data demands and drive identification. While such structures can be ported to more complex cases, this would come at the cost of increased data demands that may be simpler to list in theory than to satisfy in practice; depending on the dimensionality of the categorical variable and the underlying manifold, more complex models may or may not be practicable even with our innovations.
@@
Returning to our merchant example, consider embedding zip codes in a 2-dimensional space where the axes correspond to longitude and latitude. This approach may provide valuable insights. For example, if a retailer sells apparel suited to specific temperatures (e.g., T-shirts), projecting the high-dimensional categorical variable onto a 2-dimensional space and including each corresponding location as data features could offer a nuanced understanding of the economic problem. The model could be easily expanded by incorporating additional features, such as elevation.
Imagine, in addition to elevation, we have data on attributes like rainfall. Suppose we augmented the geographic variables with a large language model (LLM) embedding of the verbal description of each zip code. These data features may prove useful as they encompass not only the geographic characteristics of each zip code but also information on economics, culture, etc. Similarly, in our education sciences example, other data features such as average teacher salary, student-teacher ratio, and per-pupil funding may be useful for modeling the fixed effects associated with each school district or teacher education program, leveraging auxiliary information to capture the underlying characteristics of these entities and enable more accurate and efficient inference (Kane et al. 2008, Rivkin et al. 2005). Furthermore, in cases where unstructured descriptions of the category levels (e.g., school district, teacher program) are available—such as verbal descriptions of various school districts—this information can be used to augment the available structured information to further enhance estimation. This approach aligns with the use of school-level characteristics to improve the precision of estimates in education production function models, as demonstrated by Hanushek and Rivkin (1996).
@@ Conclusion
Our method has broad applicability across various research fields and industry sectors where high-dimensional categorical data is prevalent. For example, in healthcare analytics, patient data often includes categorical variables such as diagnosis codes and treatment codes. Previous work has shown the potential of using deep learning to generate unsupervised patient embeddings from electronic health records (Choi et al. 2016, Miotto et al. 2016).
Extending our approach could involve using the LLM encoding of each diagnosis code or treatment code as a set of explanatory covariates for the fixed effect. This would transform the high-dimensionality non-parametric model, which might result from including many applicable diagnosis and treatment codes, into a lower-dimensional, computationally stable, and tractable model. Such approaches may be particularly beneficial if data describing the category levels is not only structured or textual but also multi-modal, allowing for a lower-dimensional embedding to be derived from visual images, among other sources (Xu et al. 2014, Zhang et al. 2019). For instance, images of various diagnoses (e.g., an X-ray image of a fracture) could serve as inputs to the embedding to bolster its predictive accuracy. We hope our method and findings prove useful in such endeavors.